OpenStudy (spareb665):

Express ln√ab/c^3 as a sums and difference of logarithms.

OpenStudy (campbell_st):

well you will use 3 log laws to answer this question \[\ln(xy) = \ln (x) + \ln (y)\] and \[\ln(\frac{x}{y}) = \ln (x) - \ln(y) \] the last law is \[\ln(x^a) = a \ln(x)\] so I hope this helps